# Goodness-of-fit for amplitude analysis with anomaly detection

### v1

@@ 4.2 Experiment setup
We perform anomaly detection using the classifier described in Sect. 3. The pseudo-data and MC samples are randomly split into $2.5\times 10^{5}$ events for training and $2.5\times 10^{5}$ events for testing. For each signal strength case, 50 random samplings of the data are performed. For each bootstrap method, 1000 bootstrap cycles are conducted. The confidence level for all tests is set at 95%($\alpha$=0.05). Table 4 shows the rate at which each test rejects the null hypothesis that no signal is present.
@@
Figure 2 shows the empirical distributions of the p-values produced by the classifier tests for different signal strengths. The cumulative distribution functions (CDFs) in the null case are approximately uniformly distributed, indicating good control of false-positive errors. In this demonstration with the amplitude analysis of $J/\psi\to\gamma 4\pi$, it is shown that the classifier is able to detect the contribution from an additional resonance with a signal strength of 0.01 in the tests of case 1 and case 2. The detection power is sufficient for practical amplitude analyses, where contributions with fit fractions larger than 1% are typically included in the nominal fit.
Figure 2: Empirical p-value distributions under varying signal strengths for additional resonances in (a)(b) Case 1 and (c)(d) Case 2. (a)(c) Using AUC statistic; (b)(d) Using LRT statistic. The dark blue line ($\lambda=0$) represents the null case.
@@ 5 Conclusion and outlook
Multivariate goodness-of-fit tests are essential for amplitude analyses in hadron spectroscopy. We explore an ML-based anomaly detection method to evaluate goodness-of-fit for amplitude analyses of high-dimensional data. The performance study demonstrates the promise of this approach, which will be valuable for amplitude analyses of multi-body processes. Future work should incorporate systematic uncertainties when comparing amplitude analysis results to experimental data. For instance, systematic uncertainties in detection efficiency at the particle level can be addressed by reweighting the MC sample.
Acknowledgements This work is supported in part by National Natural Science Foundation of China (NSFC) under Contracts Nos.12235017, 12361141819; the Chinese Academy of Sciences Youth Team Program under Contract No. YSBR-101.

### v2

@@ 4.2 Experiment setup
We perform anomaly detection using the classifier described in Sect. 3. The pseudo-data and MC samples are randomly split into $2.5\times 10^{5}$ events for training and $2.5\times 10^{5}$ events for testing. For each signal strength case, 50 random samplings of the data are performed. For each bootstrap method, 1000 bootstrap cycles are conducted.
The p-value is calculated based on the observed test statistic and its distribution under the null hypothesis. First, the test statistic is calculated from the dataset. The null distribution that the test statistic follows under the null hypothesis is approximated using the bootstrap method. Then, the p-value can be obtained from a one-tailed hypothesis test, in which the p-value is the probability of observing a test statistic equal to or larger than the observed test statistic.
Figure 2 shows the empirical distributions of the p-values produced by the classifier tests for different signal strengths. The cumulative distribution functions (CDFs) in the null case are approximately uniformly distributed, indicating good control of false-positive errors. The empirical CDF of p-values is used to visualize the distribution of p-values obtained from a set of statistical tests. It helps in assessing whether the p-values are consistent with the null hypothesis (i.e., the two samples share the same distribution). The X-axis is p-value and the Y-axis is Cumulative Proportion, which is the fraction of p-values that are less than or equal to p. In the context of this demonstration with the amplitude analysis of $J/\psi\to\gamma 4\pi$, the power of the test is the probability that the test correctly identifies the presence of an anomaly (a deviation between the data and MC). At a confidence level of 95%($\alpha$=0.05) for all tests, Table 3 shows the rate at which each test rejects the null hypothesis that no signal is present. It is shown that the classifier is able to detect the contribution from an additional resonance with a signal strength of 0.01 in the tests of Case 1 and Case 2. The power of the test is sufficient for practical amplitude analyses, where contributions with fit fractions larger than 1% are typically included in the nominal fit.
Figure 3 shows the invariant mass distributions of $4\pi$, $3\pi$ and $2\pi$ for the data and the projects of amplitude analysis for the fit model in Case 1 with the signal strength of anomaly $\lambda=0.01$. Black dots with error bars represent ”exp” (experiment data) and the blue lines represent ”fit” (fit results). Those dashed lines represent intensity of each component in the PWA model of ”exp”. The green lines represent $J/\psi\rightarrow\gamma\eta(1760),\eta(1760)\rightarrow\rho\rho,\rho\rightarrow\pi\pi$. The brown lines represent $J/\psi\rightarrow\gamma f_{2}(2340),f_{2}(2340)\rightarrow\rho\rho,\rho\rightarrow\pi\pi$. The purple lines represent $J/\psi\rightarrow\gamma f_{0}(2100),f_{0}(2100)\rightarrow\sigma\sigma,\sigma\rightarrow\pi\pi$. The red lines represent the anomaly signal $J/\psi\rightarrow\gamma f_{0}(2100),f_{0}(2100)\rightarrow\pi a_{1}(1260),a_{1}(1260)\rightarrow\rho\pi,\rho\rightarrow\pi\pi$. The mass distribution of $3\pi$ contains two $\pi^{0}$ combinations, meaning that each event contains 2 entries in the histogram. The $\chi^{2}/N_{bin}$ value for the one-dimensional histogram is displayed on each figure. The classifier-based goodness-of-test is able to detect the anomaly, in the case that a deviation between the data and MC is difficult to be detected with the traditional $\chi^{2}$ test.
Figure 2: Empirical p-value distributions under varying signal strengths for additional resonances in (a)(b) Case 1 and (c)(d) Case 2. (a)(c) Using AUC statistic; (b)(d) Using LRT statistic. The red line ($\lambda=0$) represents the null case.
Figure 3: Invariant mass distributions of $4\pi$, $3\pi$ and $2\pi$ for the data and the projects of amplitude analysis with an additional resonance in $3\pi$ ($\lambda=0.01$). Black dots with error bars represent ”exp” (experiment data) and the blue lines represent ”fit” (fit results). Those dashed lines represent intensity of each component in the PWA model of ”exp”. The green lines represent $J/\psi\rightarrow\gamma\eta(1760),\eta(1760)\rightarrow\rho\rho,\rho\rightarrow\pi\pi$. The brown lines represent $J/\psi\rightarrow\gamma f_{2}(2340),f_{2}(2340)\rightarrow\rho\rho,\rho\rightarrow\pi\pi$. The purple lines represent $J/\psi\rightarrow\gamma f_{0}(2100),f_{0}(2100)\rightarrow\sigma\sigma,\sigma\rightarrow\pi\pi$. The red lines represent the anomaly signal $J/\psi\rightarrow\gamma f_{0}(2100),f_{0}(2100)\rightarrow\pi a_{1}(1260),a_{1}(1260)\rightarrow\rho\pi,\rho\rightarrow\pi\pi$. The mass distribution of $3\pi$ contains two $\pi^{0}$ combinations, meaning that each event contains 2 entries in the histogram.
@@ 5 Conclusion and Outlook
This study adopts a machine-learning-based approach for goodness-of-fit testing in high-dimensional amplitude analyses, addressing limitations of traditional methods like $\chi^{2}$ tests, which often fail in complex, multi-dimensional settings. By employing a probabilistic classifier (e.g., XGBoost), we have demonstrated a robust method to identify deviations between experimental data and amplitude analysis fit results. Using the $J/\psi\to\gamma 4\pi$ decay as a test case, our method successfully detected contributions from additional resonances with signal strengths as low as 1%. As an alternative to the traditional $\chi^{2}$ methods commonly used in BES and LHC experiments, our approach employs a classifier to distinguish between data and Monte Carlo samples generated from the fit model. The classifier’s output is used to construct test statistics, and the null distribution is estimated using a bootstrap method, allowing us to calculate p-values and quantitatively assess the evidence against the null hypothesis. This method provides a powerful tool for identifying deviations in high-dimensional phase spaces, enhancing the reliability of amplitude analyses. Future work will focus on several key areas. First, incorporating systematic uncertainties, such as detection efficiency and background modeling, is essential for robust goodness-of-fit assessments. Second, applying our method to other multi-body decay processes and datasets will further validate its effectiveness and general applicability. Finally, exploring advanced machine learning techniques, like deep learning models, could enhance sensitivity and accuracy in detecting subtle deviations. In summary, our machine-learning-based anomaly detection method represents a significant advancement in goodness-of-fit testing for high-dimensional amplitude analyses. By addressing key limitations of traditional methods, this approach offers a powerful tool for identifying deviations and improving the accuracy of amplitude analyses in hadron spectroscopy. Future work will refine this method and explore its broader applications in particle physics.
Acknowledgements This work is supported in part by National Natural Science Foundation of China (NSFC) under Contracts Nos.12235017, 12361141819; the Chinese Academy of Sciences Youth Team Program under Contract No. YSBR-101; the Strategic Priority Research Program of Chinese Academy of Sciences under Grant XDA0480600.